### v1

@@ 3.1 Vacuum sector
but we cannot guarantee that the averaged null energy on other semi-infinite intervals is positive; see related discussions in Davies:1976hi . It would be interesting to investigate this further.
@@
leading to the identification $\beta=2\pi\xi^{\tilde{t}}$. Physically, this says that the more the modular flow vector extends upwards into the time direction, the lower the temperature experienced by an observer following integral curves of $\xi$, and hence a corresponding decrease results in the entanglement entropy (61). We prove in appendix D that eq. (61) provides the entanglement entropy of our states $|\Psi\rangle$ obeying the condition $\tilde{b}(\tilde{x})=\tilde{c}(-\tilde{x})$, both in the vacuum and thermal sector.
@@
We then propose the following entanglement entropy formula in $d=2$ for a general state $\Psi$ with geometric modular flow $\xi$ in a subregion $\mathcal{V}=D(A)$:
@@
where $\text{d}\Sigma_{\mu}=\sqrt{|g|}\,\text{d}x\,n_{\mu}$ is the appropriate volume form on $A$, with $n^{\mu}$ the future-pointing unit normal. As a consistency check, when the flow is geometric throughout $\mathcal{V}$, $S$ is independent of the choice of Cauchy slice $A$ for $\mathcal{V}$,
@@
thanks to the conformal Killing property (8) of $\xi$ in $d=2$. Typically, $S$ will be divergent, and UV and/or IR cutoffs must be imposed on the domain of integration. In this case, for $S$ to still be conserved, the cutoff surfaces must run tangent to the modular flow $\xi$. The entropy formula (67) will be discussed further below, in section 5.
Applying eq. (67) to our excited states with geometric modular flow, we have [^29] We used $\text{d}\Sigma_{u}=-\frac{1}{2}\text{d}\Sigma^{v}=-\frac{1}{2}\sqrt{|g|}\,n^{v}\text{d}v$. Studying the line $v=-\frac{u}{\epsilon}$ as $\epsilon\to 0$, one finds $\sqrt{|g|}\,n^{v}\to 1$.
@@
Consider then our excited states. Their vacuum-subtracted entanglement entropy is
@@
We close this subsection with a consistency check of the entropy formula (67): We test that it yields a relative entropy with the expected properties. The relative entropy of $\rho_{\Psi}$ with respect to $\rho_{\Omega}$ is defined as
@@ 3.4 Disentangled states
We can repeat the same procedure for the Borchers flow (36):
@@
At small $\tilde{u}$, this agrees with the disentangled vacuum state result (114). At large $\tilde{u}$, we obtain $\frac{c\pi}{16\beta^{2}}$, which is $\frac{3}{4}$ of $\frac{c\pi}{12\beta^{2}}$, the constant thermal state null energy density in eq. (83).
@@ 4.1 Vacuum flows
Moreover, since $(u_{0},v_{0})$ is spacelike separated from $\mathcal{D}$, while $(u_{s},v_{s})$ is timelike separated from $\mathcal{D}$ for $s>s_{0}$, we conclude eq. (124) can send spacelike separated points to timelike separated points. This is a special feature of Minkowski CFTs and does not arise when considering CFTs on the cylinder. As discussed in appendix E, making this effect consistent with microcausality requires a modification of the usual, local transformation rule eq. (15) for primary operators. While operators transform in the usual local way at finite distances in the spacetime, they transform nonlocally upon crossing $\mathcal{I}^{\pm}$.
@@ 4.2 A state for every flow in the diamond and in the cylinder
In footnote 16, we remarked that SCTs are symmetries of the vacuum sector. Section 2 further argued that states within the same sector share similar asymptotic properties, as GNS Hilbert spaces are generated by locally smeared quantum fields. This appears to be in tension with the multiplicative ambiguity in the modular flow falloffs (133), arising from the use of SCTs to bring infinity to a finite point. The resolution lies in the energy density formula (88). In the Rindler frame, rescaling modular flow falloffs (13) alters the energy density falloffs from $\langle T_{tt}\rangle\sim\frac{1}{x^{4}}$ to $\langle T_{tt}\rangle\sim\frac{1}{x^{2}}$. Contrastingly, in the diamond frame, the modular flow rescalings in eq. (131) maintain the decay $\langle T_{tt}\rangle\sim\frac{1}{x^{4}}$. Therefore, if vacuum-sector states are identified by the asymptotic behavior of the energy density, there is no tension; we should allow for the multiplicative ambiguity in the diamond frame because it leaves the energy decay invariant. We implicitly employed this reasoning in eq. (41).
@@ 5.1 Entanglement currents
Accordingly, we obtain a generalization of eq. (67) for $d>2$,
@@
$S=\alpha_{0}\int\sqrt{g}\,\text{d}x\,\text{d}\Omega_{d-2}=\alpha_{0}\,R^{d-1}\,\Omega_{d-2}\int^{u_{\max}}_{0}\text{d}u\,\sinh^{d-2}\!u\,,$ (148)
where $g$ is the determinant of the $H^{d-1}$ metric and $\Omega_{d-2}=\frac{2\pi^{\frac{d-1}{2}}}{\Gamma\left(\frac{d-1}{2}\right)}$ is the area of the unit $(d-2)$-sphere. Moreover, $\alpha_{0}$ is a thermal entropy density; c.p. eqs. (138)-(140) with $\xi=R\partial_{\tau}$ from footnote 45. It can be written as
@@ 6 Discussion
We can interpret the entropy current as a book-keeping device tracking the amount of entanglement we must add to obtain the state of interest from a corresponding disentangled state. Heuristically, we can think of adding a CFT plasma with a local temperature which varies from point to point. This temperature may even change from one moment to the next along the worldlines of observers following the modular flow. The conservation of $j_{S}$ remarkably indicates that these changes in temperature precisely balance among the collection of observers such that the total entanglement entropy is unchanged with any choice of Cauchy slice for the subregion of interest.
@@
Nevertheless, it would be interesting to see whether ideas from geometric modular flows may be useful in the study of entanglement contours. In particular, local temperatures can be defined even in states whose modular Hamiltonian is not of the form (5) Arias:2016nip . It would be therefore interesting to understand whether a temperature-based entanglement current such as eq. (139) can be defined even in these states, providing a general realization of an entanglement density, or contour. Indeed, this is suggested by the results of appendix F, where we show that eq. (140) gives the correct entanglement entropy for a state whose modular Hamiltonian contains nonlocal terms.
@@
In sections 2 to 4, we built excited states with geometric modular flows in $d=2$ starting from either the vacuum state or the thermal state. It is natural to ask whether we can build states with geometric modular flows starting instead from primary states.
Consider a local primary operator $V(x)$ acting on the vacuum state at the origin of a Euclidean CFT2 on the plane,
@@
Under radial quantization, this defines a state on circles centered at the origin. With a suitable conformal mapping and Wick rotation, this gives a state on the Lorentzian cylinder.
In holography, an example of a primary state on the Lorentzian cylinder is the dual of the AdS3 conical defect geometry Deser:1983nh ; Deser:1983tn ; Banerjee:2016qca . This geometry is obtained from empty AdS3 by reducing the periodicity of the spatial $\varphi$-circle from $2\pi$ to $2\pi\lambda$, with $\lambda<1$. In this case, it can be argued using entanglement wedge reconstruction Dong:2016eik that the modular flow associated with small causal diamonds in the CFT, whose entanglement wedge does not contain the defect, is a simple rescaling of the vacuum modular flow with $t\to\frac{t}{\lambda}$ and $\varphi\to\frac{\varphi}{\lambda}$. On the other hand, the modular flow is non-geometric for a boundary subregion which is so large that its entanglement wedge contains the defect. Our result on the existence of a state for any one-sided Unruh flow straightforwardly carry over to the case of the small causal diamonds $\mathcal{D}_{c}$ in such primary states. That is, a $\mathcal{D}_{c}$-preserving conformal transformation $U_{\varphi}$ maps $|V\rangle$ into a new state $U_{\varphi}|V\rangle$ with a different geometric modular flow in $\mathcal{D}_{c}$. On the other hand, the flow will be non-geometric in the complementary diamond. It may be interesting to study how the modular flow changes as the size of $\mathcal{D}_{c}$ increases in this setting.
@@
Two sectors at different temperatures are unitarily inequivalent; see footnote 33. Consistency with this statement, in our framework, is imposed when requiring that the conformal unitaries act trivially at spatial infinity, such that the vacuum or thermal sector falloffs for the modular flow in eqs. (13) and (40) are preserved. One may wonder how to extend our construction to allow for operators which change the temperature. In this setting, it is important to understand the asymptotic behavior which characterizes the different sectors.
@@
precisely matches the Unruh temperature measured by the observer’s detector. However, this is a special case and in general $\beta=2\pi\norm{\xi}$ is not necessarily identified with the physical inverse temperature that an observer following the modular flow experiences.
To illustrate this point, we consider the example of the vacuum flow (123) in a causal diamond. The integral curves corresponding to this flow can be obtained as in eq. (124),
@@
where we specify the curves by where they cross the $x$-axis (i.e., $t=0$) with $v_{0}=x_{0}=-u_{0}$, which occurs at $s=0$. Now consider these curves (177) to be the worldlines describing a family of Unruh observers traversing the diamond. Hence, we may wish to re-express them in terms of the proper time $\tau$ of the observers. This is done using the relation [^56] To derive this, integrate the relation $\frac{\text{d}\tau}{\text{d}s}=\norm{\xi}$ from footnote 28, using eqs. (177) and (153). We have chosen the integration constant such that $\tau=0$ corresponds to $s=0$. Note that in traversing the entire diamond, $-\infty<s<\infty$ while the proper time $\tau$ spans a finite range $-\tau_{\rm max}<\tau<\tau_{\rm max}$ where at its maximum value one has $\cosh\frac{x_{0}\,\tau_{\rm max}}{R^{2}-x_{0}^{2}}=\frac{R}{\sqrt{R^{2}-x_{0}^{2}}}$.
@@
Now given the periodicity in eq. (179) and the symmetry $\Delta s(\tau_{1},\tau_{2})=\Delta s(\tau_{2},\tau_{1})$, the Wightman function then satisfies $W(\tau_{1}+i\beta_{\rm phys},\tau_{2})=W(\tau_{2},\tau_{1})$, which implies a thermal response of the detector at temperature $\beta_{\mathrm{phys}}$, as shown rigorously in Birrell:1982ix ; DeBievre:2006pys . That is, the observer experiences a thermal bath with temperature given by [^58] We could arrive at the same conclusion by noting that the observers described by eq. (177) are all uniformly accelerating, with $a(x_{0})=\frac{2x_{0}}{R^{2}-x_{0}^{2}}$. Hence, eq. (181) has the same form as the standard Unruh-Bisognano-Wichmann temperature, $T=\frac{a(x_{0})}{2\pi}$. Similar statements apply for the geometric flow in the causal complement of the diamond, since the observers there also follow trajectories with constant acceleration. , [^59] One may be concerned that the detector will not reach thermal equilibrium because our observers traverse the causal diamond in a finite time — see footnote 56. But the observer can simply initialize the detector with its different energy levels populated in a thermal distribution determined by $\beta_{\rm phys}$. Then, apart from small fluctuations, the observer’s detector remains in equilibrium as she crosses the diamond.
@@
A few comments are in order here. For $x_{0}=0$, eq. (178) reduces to $\tanh{\frac{s}{2}}=\frac{\tau}{R}$ and the corresponding trajectory (177) becomes $u(\tau)=v(\tau)=\tau$, hence describing a stationary observer with constant $x$-coordinate. In this case, $T_{\mathrm{phys}}$ vanishes, as to be expected for an inertial observer in the Minkowski vacuum state. On the other hand, as $x_{0}$ approaches the entangling surface at $x_{0}=\pm R$, the modular flow approaches a boost flow and $T_{\mathrm{phys}}$ approaches the standard Unruh-Bisognano-Wichmann temperature $T\simeq\frac{1}{2\pi\,\delta x}$, where $\delta x=R-|x_{0}|$ is the distance to the entangling surface. Further, note that identical results hold for the double-cone region in $d>2$, since the vacuum modular flow (153) for the double-cone takes the same form in lightcone coordinates in all dimensions. Similarly, identical results hold in any theory where the corresponding Wightman functions depend only on the proper distance.
We now compare the physical temperature $T_{\mathrm{phys}}$ experienced by observers following the diamond flow with the modular temperature appearing in our entanglement entropy formula, $\beta=2\pi\norm{\xi}$. From eq. (123), $\beta=2\pi\norm{\xi}=\frac{\pi}{R}\sqrt{(R^{2}-u^{2})(R^{2}-v^{2})}$, which yields
@@
This does not match the physical temperature in eq. (181). Indeed, it assigns a nonzero temperature for the inertial observer with $x_{0}=0$! From this expression, one may also be concerned that the modular temperature will become negative for large values of $\tau$. However, the observers cross the causal diamond in finite proper time — see footnote 56. A more detailed examination shows that $T$ diverges if we follow the observers to either the future or past tips of the causal diamond, but $T$ never becomes negative.
Recall, the modular temperature can also be written as $\beta=2\pi\frac{\text{d}\tau}{\text{d}s}$ (see footnote 28), which is the temperature ascribed to observers in Martinetti:2002sz . [^60] See related discussions in Earman:2011zz . While the authors do not associate a physical meaning to this temperature there, they propose that it should be related to the fact that observers do not access all the degrees of freedom of the underlying quantum field theory. We see this intuition is correct in that the modular temperature provides a simple prescription (140) to evaluate the entanglement entropy for region which the observers are able to access.
@@
where the final equality holds when $\xi$ is Killing. In this case, we then have $\beta=2\pi\frac{\text{d}\tau}{\text{d}s}=\beta_{\rm phys}$. That is, the modular temperature and the physical temperature experienced by the observer coincide, as with the vacuum Rindler flow. Conversely, when the geometric modular flow is described by a general conformal Killing vector, as is the case for the flow (124) on the causal diamond, the divergence in eq. (183) is nonvanishing and hence $\beta\neq\beta_{\mathrm{phys}}$. [^61] In changing conformal frames, as in section 4, a conformal Killing vector may be transformed into an ordinary Killing field. This happened in section 5.1, when conformally mapping the double cone to a hyperbolic spacetime, and below figure 10, when conformally mapping the Milne wedge to a hyperbolic spacetime. In these situations thermal effects become physical, but only in a particular conformal frame.

### v2

@@ 3.1 Vacuum sector
but we cannot guarantee that the averaged null energy on other semi-infinite intervals is positive; see related discussions in Davies:1976hi .
@@
leading to the identification $\beta=2\pi\xi^{\tilde{t}}$. Physically, this says that the more the modular flow vector extends upwards into the time direction, the lower the temperature associated with an observer following integral curves of $\xi$, and hence a corresponding decrease results in the entanglement entropy (61). We prove in appendix D that eq. (61) provides the entanglement entropy of our states $|\Psi\rangle$ obeying the condition $\tilde{b}(\tilde{x})=\tilde{c}(-\tilde{x})$, both in the vacuum and thermal sector.
@@
We accordingly propose the following entanglement entropy formula in $d=2$ for a general state $\Psi$ with geometric modular flow $\xi$ in a subregion $\mathcal{V}=D(A)$:
@@
where $\text{d}\Sigma_{\mu}=\sqrt{h}\,\text{d}x\,n_{\mu}$ is the appropriate volume form on $A$, with $n^{\mu}$ the future-pointing unit normal. As a consistency check, when the flow is geometric throughout $\mathcal{V}$, $S$ is independent of the choice of Cauchy slice $A$ for $\mathcal{V}$,
@@
thanks to the conformal Killing property (8) of $\xi$ in $d=2$. Typically, $S$ will be divergent, and UV and/or IR cutoffs must be imposed on the domain of integration. In this case, for $S$ to still be conserved, the cutoff surfaces must run tangent to the modular flow $\xi$. The entropy formula (67) will be discussed further in section 5 and proved in section 6.
Applying eq. (67) to our excited states with geometric modular flow, we have [^30] We used $\text{d}\Sigma_{u}=-\frac{1}{2}\text{d}\Sigma^{v}=-\frac{1}{2}\sqrt{h}\,n^{v}\text{d}v$. Studying the line $v=-\frac{u}{\epsilon}$ as $\epsilon\to 0$, one finds $\sqrt{h}\,n^{v}\to 1$.
@@
Now consider our excited states. Their vacuum-subtracted entanglement entropy is
@@
We close this subsection with a consistency check of the entropy formula (67): We test that it yields a relative entropy with the expected properties. [^32] For an analysis of the relative entropy of CFT2 states excited by a conformal transformation, see deBoer:2023lrd . The relative entropy of $\rho_{\Psi}$ with respect to $\rho_{\Omega}$ is defined as
@@ 3.4 Disentangled states
Note, eq. (114) implies the energy density $\langle T_{\tilde{t}\tilde{t}}\rangle_{\Omega_{\mathrm{dis}}}$ is negative for all nonzero $\tilde{u}$ and $\tilde{v}$. On the other hand, we expect the total energy of |$\ket{\Omega_{\mathrm{dis}}}$ is positive, because we can approximate |$\ket{\Omega_{\mathrm{dis}}}$ using a sequence of smooth conformal maps, each of which yields a positive total energy according to eq. (53). In practice, the positive total energy arises due to a singular, positive lump of energy density localized at the entangling surface, as explained in Parentani:1993yz for the Rindler vacuum.
We now repeat the disentangling procedure for the Borchers flow (36):
@@
At small $\tilde{u}$, this agrees with the disentangled vacuum state result (114). At large $\tilde{u}$, we obtain $\frac{c\pi}{16\beta^{2}}$, which is $\frac{3}{4}$ of $\frac{c\pi}{12\beta^{2}}$, the constant thermal state null energy density in eq. (83). As with the Rindler vacuum, we expect the total energy here is positive.
@@ 4.1 Vacuum flows
[…]
@@ 4.2 A state for every flow in the diamond and in the cylinder
In footnote 17, we remarked that SCTs are symmetries of the vacuum sector. Section 2 further argued that states within the same sector share similar asymptotic properties, as GNS Hilbert spaces are generated by locally smeared quantum fields. This appears to be in tension with the multiplicative ambiguity in the modular flow falloffs (133), arising from the use of SCTs to bring infinity to a finite point. The resolution lies in the energy density formula (88). In the Rindler frame, rescaling modular flow falloffs (13) alters the energy density falloffs from $\langle T_{tt}\rangle\sim\frac{1}{x^{4}}$ to $\langle T_{tt}\rangle\sim\frac{1}{x^{2}}$. Contrastingly, in the diamond frame, the modular flow rescalings in eq. (131) maintain the decay $\langle T_{tt}\rangle\sim\frac{1}{x^{4}}$. Therefore, if vacuum-sector states are identified by the asymptotic behavior of the energy density, there is no tension; we should allow for the multiplicative ambiguity in the diamond frame because it leaves the energy decay invariant. We implicitly employed this reasoning in eq. (41).
@@ 5.1 Entanglement currents
Accordingly, we obtain a generalization of eq. (67) for $d>2$, [^47] This formula is Weyl invariant. To see why, write $S=-\alpha\int_{A}\sqrt{h}\,\text{d}^{d-1}x\frac{n^{\nu}}{\norm{n}}g_{\mu\nu}\frac{\xi^{\mu}}{\norm{\xi}^{d}}$ and use that under a Weyl transformation, $g_{\mu\nu}\to\Omega^{2}g_{\mu\nu}$, $\sqrt{h}\to\Omega^{d-1}\sqrt{h}$, $\norm{n}\to\Omega\norm{n}$, and $\norm{\xi}\to\Omega\norm{\xi}$.
@@
$S=\alpha_{0}\int\sqrt{h}\,\text{d}x\,\text{d}\Omega_{d-2}=\alpha_{0}\,R^{d-1}\,\Omega_{d-2}\int^{u_{\max}}_{0}\text{d}u\,\sinh^{d-2}\!u\,,$ (148)
where $h$ is the determinant of the $H^{d-1}$ metric and $\Omega_{d-2}=\frac{2\pi^{\frac{d-1}{2}}}{\Gamma\left(\frac{d-1}{2}\right)}$ is the area of the unit $(d-2)$-sphere. Moreover, $\alpha_{0}$ is a thermal entropy density; c.p. eqs. (138)-(140) with $\xi=R\partial_{\tau}$ from footnote 49. It can be written as
@@ 6 Discussion
Multiple subregions
So far, we have studied modular flows of connected regions, e.g., the Rindler wedge or a causal diamond. A generalization of our construction comes from considering subregions with multiple disconnected components.
Let $\mathcal{V}$ describe two spacelike-separated, finite regions, $\mathcal{V}=\mathcal{V}_{1}\cup\mathcal{V}_{2}$, with a finite separation. For a typical state, the modular Hamiltonian with respect to $\mathcal{V}$ contains nonlocal terms mixing $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ which represent entanglement between the two subregions (for example, see eq. (279) in appendix F). However, we expect there exist special states for which $\mathcal{V}=\mathcal{V}_{1}\cup\mathcal{V}_{2}$ and yet the modular Hamiltonian is local in $\mathcal{V}$, leading to vanishing entanglement between $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$.
This is related to the following general statement Witten:2018zxz ; Fewster:2019ixc ; Roos:1970fm ; Buchholz:1973bk ; Doplicher:1984zz : For well-behaved QFTs in Minkowski spacetime, given any two states |$\ket{\Psi_{1}}$ and |$\ket{\Psi_{2}}$ in a GNS Hilbert space $\mathcal{H}$, there exists a (non-unique) state |$\ket{\Psi_{p}}\in\mathcal{H}$ which is indistinguishable from |$\ket{\Psi_{1}}$ for measurements in $\mathcal{V}_{1}$ and indistinguishable from |$\ket{\Psi_{2}}$ for measurements in $\mathcal{V}_{2}$. That is, for ${\sf a}_{1}\in\mathcal{A}_{\mathcal{V}_{1}}$ and ${\sf a}_{2}\in\mathcal{A}_{\mathcal{V}_{2}}$,
$\langle\Psi_{p}|{\sf a}_{1}|\Psi_{p}\rangle=\langle\Psi_{1}|{\sf a}_{1}|\Psi_{1}\rangle\qquad\langle\Psi_{p}|{\sf a}_{2}|\Psi_{p}\rangle=\langle\Psi_{2}|{\sf a}_{2}|\Psi_{2}\rangle\,.$ (175)
Further, one can choose $|\Psi_{p}\rangle$ to be a product state, meaning
$\langle\Psi_{p}|{\sf a}_{1}{\sf a}_{2}|\Psi_{p}\rangle=\langle\Psi_{1}|{\sf a}_{1}|\Psi_{1}\rangle\langle\Psi_{2}|{\sf a}_{2}|\Psi_{2}\rangle$ (176)
and there is no entanglement between $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ in the state $|\Psi_{p}\rangle$. From Takesaki:ToOA , the modular flow $U_{\Psi_{p}}^{\mathcal{V}}(s)$ then factorizes, i.e. it acts as $U_{\Psi_{1}}^{\mathcal{V}_{1}}(s)$ on all operators in the $\mathcal{V}_{1}$ algebra, and $U_{\Psi_{2}}^{\mathcal{V}_{2}}(s)$ on all operators in the $\mathcal{V}_{2}$ algebra.
To see the relation with local modular Hamiltonians, consider the case $d=2$ and Unruh flows $\xi_{1}$ in $\mathcal{V}_{1}$ and $\xi_{2}$ in $\mathcal{V}_{2}$. By our results in section 2, there exist vacuum-sector CFT2 states $|\Psi_{1}\rangle$ and $|\Psi_{2}\rangle$ whose modular Hamiltonians generate $\xi_{1}$ in $\mathcal{V}_{1}$ and $\xi_{2}$ in $\mathcal{V}_{2}$, respectively. Now using the result above, there exists a vacuum-sector state |$\ket{\Psi_{p}}$ whose modular flow with respect to $\mathcal{V}$ factorizes, generating the local flows $\xi_{1}$ in $\mathcal{V}_{1}$ and $\xi_{2}$ in $\mathcal{V}_{2}$.
In summary, we can specify an arbitrary Unruh flow $\xi_{1}$ in $\mathcal{V}_{1}$ and an arbitrary Unruh flow $\xi_{2}$ in $\mathcal{V}_{2}$, and we expect there exists a state in the CFT2 vacuum sector with the corresponding modular flow. Hence, our result on the existence of a state for every Unruh flow in Minkowski CFT2 generalizes to multi-component subregions, and the states of interest are disentangled between the various subregions. [^58] Unlike in section 3.4, there are no singularities associated with the disentangled state. This is because the finite separation between $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ ensures the quantum correlations are UV-finite.
We apply this result to a particular example in appendix F. There, we consider free fermion theory in $d=2$. In this case, the one-sided modular Hamiltonian $h_{\Omega}^{\mathcal{V}}$ for the vacuum state with respect to $\mathcal{V}_{1}\cup\mathcal{V}_{2}$ is comprised of a local term as well as a nonlocal term mixing the two subregions Casini:2016fgb . Using the discussion above, there exists a state |$\ket{\Psi_{p}}$ whose modular Hamiltonian is given by the local terms in $h_{\Omega}^{\mathcal{V}}$. This is useful because we expect our entanglement entropy formula eq. (140) applies only when the modular Hamiltonian is local, hence for the subregion $\mathcal{V}$ it can be applied to |$\ket{\Psi_{p}}$ but not |$\ket{\Omega}$. In appendix F, we calculate the entanglement entropy of $\mathcal{V}$ in the state |$\ket{\Psi_{p}}$ and show that it matches that of the vacuum state |$\ket{\Omega}$, calculated in Casini:2016fgb . This implies that the nonlocal terms in $h_{\Omega}^{\mathcal{V}}$ do not actually contribute to the entanglement entropy, reinforcing the interpretation of the nonlocal terms as representing entanglement between $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$, rather than entanglement between $\mathcal{V}$ and its complement. Our analysis in appendix F shows that this result easily extends to the case of $n>2$ subregions.
Having generalized our analysis of Unruh modular flows to the case of multi-component subregions, resulting in local expressions for certain corresponding entanglement entropies, we now explore further the relationship between local entropy densities and entanglement structures.
@@
We can interpret the entropy current as a book-keeping device tracking the amount of entanglement we must add to obtain the state of interest from a corresponding disentangled state. Heuristically, we can think of adding a CFT plasma with a local temperature which varies from point to point. This temperature may even change from one moment to the next along the worldlines of observers following the modular flow. The conservation of $j_{S}$ indicates that these changes in temperature precisely balance among the collection of observers such that the total entanglement entropy is unchanged with any choice of Cauchy slice for the subregion of interest.
@@
Nevertheless, it would be interesting to see whether ideas from geometric modular flows may be useful in the study of entanglement contours. In particular, as discussed further below, local temperatures can be defined even in states whose modular Hamiltonian is not of the form (5) Arias:2016nip . It would therefore be interesting to understand whether a temperature-based entanglement current such as eq. (139) can be defined even in these states, providing a general realization of an entanglement density or contour. Indeed, this is suggested by the results of appendix F, where we show that eq. (140) gives the correct entanglement entropy for a state whose modular Hamiltonian contains not only local “temperature” terms but also additional nonlocal terms.
@@
In sections 2 to 4, we built excited states with geometric modular flows with respect to connected subregions, starting from either the vacuum state or the thermal state. In many cases, it is simple to generalize this construction to the case of primary states and/or multi-component subregions.
A primary state can be defined by acting with a local primary operator $V(x)$ on the vacuum state at the origin of a Euclidean CFT2 on the plane,
@@
Under radial quantization, this defines a state on circles centered at the origin. With a suitable conformal mapping and Wick rotation, it defines a state on the Lorentzian cylinder.
In holography, an example of a primary state on the Lorentzian cylinder is the dual of the AdS3 conical defect geometry Deser:1983nh ; Deser:1983tn ; Banerjee:2016qca . This geometry is obtained from empty AdS3 by reducing the periodicity of the spatial $\varphi$-circle from $2\pi$ to $2\pi\lambda$, with $\lambda<1$. In this case, it can be argued using entanglement wedge reconstruction Dong:2016eik that the modular flow associated with small causal diamonds in the CFT, whose entanglement wedge does not contain the defect, is a simple rescaling of the vacuum modular flow with $t\to\frac{t}{\lambda}$ and $\varphi\to\frac{\varphi}{\lambda}$. On the other hand, the modular flow is non-geometric for a boundary subregion which is so large that its entanglement wedge contains the defect. Our result on the existence of a state for any one-sided Unruh flow straightforwardly carries over to the case of the small causal diamonds $\mathcal{D}_{c}$ in such primary states. That is, a $\mathcal{D}_{c}$-preserving conformal transformation $U_{\varphi}$ maps $|V\rangle$ into a new state $U_{\varphi}|V\rangle$ with a different geometric modular flow in $\mathcal{D}_{c}$. Of course, the flow will remain non-geometric in the complementary diamond. It may be interesting to study how the modular flow changes as the size of $\mathcal{D}_{c}$ increases in this setting.
@@
Two sectors at different temperatures are unitarily inequivalent; see footnote 35. Consistency with this statement, in our framework, follows from the requirement that the conformal unitaries act trivially at spatial infinity, such that the vacuum or thermal sector falloffs for the modular flow in eqs. (13) and (40) are preserved. One may wonder how to extend our construction to allow for operators which change the temperature. In this setting, it is important to understand the asymptotic behavior which characterizes the different sectors.
@@
precisely matches the physical temperature measured by the observer’s detector. However, this is a special case; in general, $\beta=2\pi\norm{\xi}$ is not necessarily identified with the physical inverse temperature that an observer following the modular flow experiences.
To illustrate this point, consider the example of the vacuum flow (123) in a causal diamond. The integral curves corresponding to this flow can be obtained as in eq. (124),
@@
where we specify the curves by where they cross the $x$-axis (i.e., $t=0$) with $v_{0}=x_{0}=-u_{0}$, which occurs at $s=0$. Now consider these curves (179) to be the worldlines describing a family of Unruh observers traversing the diamond. Hence, we may wish to re-express them in terms of the proper time $\tau$ of the observers. This is done using the relation [^61] To derive this, integrate the relation $\frac{\text{d}\tau}{\text{d}s}=\norm{\xi}$ from footnote 29, using eqs. (179) and (153). We have chosen the integration constant such that $\tau=0$ corresponds to $s=0$. Note that in traversing the entire diamond, $-\infty<s<\infty$ while the proper time $\tau$ spans a finite range $-\tau_{\rm max}<\tau<\tau_{\rm max}$ where at its maximum value one has $\cosh\frac{x_{0}\,\tau_{\rm max}}{R^{2}-x_{0}^{2}}=\frac{R}{\sqrt{R^{2}-x_{0}^{2}}}$.
@@
Now given the periodicity in eq. (181) and the symmetry $\Delta s(\tau_{1},\tau_{2})=\Delta s(\tau_{2},\tau_{1})$, the Wightman function then satisfies $W(\tau_{1}+i\beta_{\rm phys},\tau_{2})=W(\tau_{2},\tau_{1})$, which implies a thermal response of the detector at temperature $\beta_{\mathrm{phys}}$, as shown rigorously in Birrell:1982ix ; DeBievre:2006pys . That is, the observer experiences a thermal bath with temperature given by [^63] One may be concerned that the detector will not reach thermal equilibrium because our observers traverse the causal diamond in a finite time — see footnote 61. But the observer can simply initialize the detector with its different energy levels populated in a thermal distribution determined by $\beta_{\rm phys}$. Then, apart from small fluctuations, the observer’s detector remains in equilibrium as she crosses the diamond.
@@
A few comments are in order here. For $x_{0}=0$, eq. (180) reduces to $\tanh{\frac{s}{2}}=\frac{\tau}{R}$ and the corresponding trajectory (179) becomes $u(\tau)=v(\tau)=\tau$, hence describing a stationary observer with constant $x$-coordinate. In this case, $T_{\mathrm{phys}}$ vanishes, as to be expected for an inertial observer in the Minkowski vacuum state. On the other hand, as $x_{0}$ approaches the entangling surface at $x_{0}=\pm R$, the modular flow approaches a boost flow and $T_{\mathrm{phys}}$ approaches the standard Unruh-Bisognano-Wichmann temperature $T\simeq\frac{1}{2\pi\,\delta x}$, where $\delta x=R-|x_{0}|$ is the distance to the entangling surface. More generally, we observe that the vacuum flow in the causal diamond has a special feature: the observers are following trajectories of constant acceleration $a(x_{0})=2x_{0}/(R^{2}-x_{0}^{2})$. Hence, eq. (183) has the form of the standard Unruh-Bisognano-Wichmann temperature, $T=\frac{a}{2\pi}$. [^64] Similar statements apply for the geometric flow in the causal complement of the diamond, since the observers there also follow trajectories with constant acceleration. Identical results hold for the double-cone region in $d>2$, since the vacuum modular flow (153) for the double-cone takes the same form in lightcone coordinates in all dimensions. Further, identical results hold in any theory where the corresponding Wightman functions depend only on the proper distance.
We now contrast the physical temperature $T_{\mathrm{phys}}$ experienced by observers following the diamond flow with the modular temperature $\beta=2\pi\norm{\xi}$ appearing in our entanglement entropy formula. From eq. (123), $\beta=2\pi\norm{\xi}=\frac{\pi}{R}\sqrt{(R^{2}-u^{2})(R^{2}-v^{2})}$, which yields
@@
This does not match the physical temperature in eq. (183). Indeed, it assigns a nonzero temperature for the inertial observer with $x_{0}=0$! From this expression, one may also be concerned that the modular temperature will become negative for large values of $\tau$. However, the observers cross the causal diamond in finite proper time — see footnote 61. A more detailed examination shows that $T$ diverges if we follow the observers to either the future or past tips of the causal diamond, but $T$ never becomes negative.
Recall, the modular temperature can also be written as $\beta=2\pi\frac{\text{d}\tau}{\text{d}s}$ (see footnote 29), which is the temperature ascribed to observers in Martinetti:2002sz . [^65] See related discussions in Earman:2011zz . While the authors do not associate a physical meaning to this temperature, they propose it should be related to the fact that observers do not access all the degrees of freedom of the underlying quantum field theory. We see this intuition is correct in that the modular temperature provides a simple prescription (140) to evaluate the entanglement entropy for the region the observers can access.
@@
where the final equality holds when $\xi$ is Killing. In this case, we then have $\beta=2\pi\frac{\text{d}\tau}{\text{d}s}=\beta_{\rm phys}$. That is, the modular temperature and the physical temperature experienced by the observer coincide, as with the vacuum Rindler flow. Conversely, when the geometric modular flow is described by a general conformal Killing vector, as is the case for the flow (124) on the causal diamond, the divergence in eq. (185) is nonvanishing and hence $\beta\neq\beta_{\mathrm{phys}}$.